# Towards the understanding of heavy quarks hadronization: from leptonic to heavy-ion collisions

### v2

@@ III.3 Proton–proton collisions
The low-$p_{\mathrm{T}}$ charm baryon enhancement is reflected in the $p_{\mathrm{T}}$-integrated yields and their ratios to the $\mathrm{D^{0}}$ meson. The left panel of Fig. 15 shows for pp collisions at $\sqrt{s}=5.02~{}\mathrm{TeV}$ the $p_{\mathrm{T}}$-integrated production cross sections at midrapidity per unit of rapidity for various open- and hidden-charm mesons ($\mathrm{D^{+}}$, $\mathrm{D_{s}^{+}}$, $\mathrm{D^{*+}}$, and $\mathrm{J}/\psi$[188, 157, 189] and baryons ($\Lambda_{\mathrm{c}}^{+}$ and $\Xi_{\mathrm{c}}^{0}$[77, 179], taking the average of particles and antiparticles, and normalizing to the $\mathrm{D^{0}}$ meson. The experimental data is compared with results of pp simulations done with PYTHIA 8 using both the Monash tune [11] and the tunes implementing CR-BLC [33]. For the open charm meson ratios, the PYTHIA 8 predictions obtained with the different tunes are fairly similar and describe the data within uncertainty. The value of the $\mathrm{D^{*+}}/\mathrm{D^{0}}$ ratio obtained with Monash2013 setting $StringFlav:mesonCvector=1.5$ is also reported. As anticipated in Sect. II.1, values of this parameter around 1.5 or larger are preferable to better describe the data, as shown in the right panel of the same figure for the $p_{\mathrm{T}}$-differential $\mathrm{D^{*+}}$/$\mathrm{D^{0}}$ ratio. However, this is not an artefact of the modelling of $D^{*+}$ pp collision environments. A discrepancy is existent also when considering $\mathrm{e^{+}e^{-}}$ collisions data from LEP and is a consequence of the parameters StringFlav:mesonCvector and correspondingly StringZ:rFactC being mistuned. The default value of 0.88 for StringFlav:mesonCvector should instead be closer to $\sim$1.5. Contrary to the meson case, and as expected from what discussed about the low-$p_{\mathrm{T}}$ trend of the baryon-to-meson ratios shown in Fig. 13, significant differences between the model expectations and data are observed when considering the baryon-to-$\mathrm{D^{0}}$ ratios. The Monash 2013 tune is observed to underestimate the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ and $\Xi_{\mathrm{c}}^{0}/\mathrm{D^{0}}$ ratios by nearly $8\,\sigma$ and $2.3\,\sigma$, respectively. Comparatively, the CR tunes provide larger baryon-to-meson ratios with the Mode 2 tune well reproducing the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio. Nonetheless, they still underestimate the $\Xi_{\mathrm{c}}^{0}/\mathrm{D^{0}}$ ratio by about $2\,\sigma$, which again is unsurprising when looking at the $p_{\mathrm{T}}$ distributions in Fig. 13. For the $\mathrm{J}/\psi/\mathrm{D^{0}}$ ratio the CR tunes provide a better description than the Monash tune, however, all PYTHIA 8 tunes underestimate the measurement. One should note that in the simulations, as in the experimental measurement, the $\mathrm{J}/\psi$ cross section consists of the prompt and beauty feed-down contributions. From these results, one can conclude that the high baryon-to-meson ratios ($\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$, $\Sigma_{\mathrm{c}}^{0,+,++}$/$\mathrm{D^{0}}$, $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$, and $\Omega_{\mathrm{c}}^{0}$/$\mathrm{D^{0}}$), with magnitudes much larger than those of FF ratio measured in $\mathrm{e^{+}e^{-}}$ collisions, clearly indicate that the hadronisation process is not universal and that modelling it solely with FF and FFunc determined from $\mathrm{e^{+}e^{-}}$ data fails already in pp collisions. The right panel of Fig. 15 shows the same experimental data compared in addition to the SHM+RQM and Catania models. For the SHM+RQM for a hadronisation temperature $T_{\mathrm{h}}=$ 170 MeV. Similarly to what is observed for the $p_{\mathrm{T}}$ differential ratios, the SHM+RQM model describes the $\Lambda_{\mathrm{c}}^{+}$ measurements but underestimates the $\Xi_{\mathrm{c}}^{0}$ one. On the other hand, the Catania model overpredicts the production of the $\Lambda_{\mathrm{c}}^{+}$ baryon but captures the production of the $\Xi_{\mathrm{c}}^{0}$ baryon. For the Catania model the $D^{*+}/D^{0}$ value shown is the one for the mesons coming from coalescence, because the fragmentation functions employed are inclusive and give only the total contribution to $D$ meson including the $D^{*}$ feed-down. The significantly large $\Lambda_{\mathrm{c}}^{+}/D^{0}$ is due to the combination of both a lower $D^{0}$ and a larger $\Lambda_{\mathrm{c}}^{+}$ yield with respect to experimental data. However, it has to be noticed that the value in the plot is the original prediction employing the width corresponding to the mean square radii of hadrons from the quark model, hence without any free parameter; a moderate modification of the width parameter would allow to get quite close to the data.
@@
Figure 17: Left panel: ratios $\mathrm{\Lambda_{\mathrm{b}}^{0}}/\mathrm{B^{0}}$ as a function of $p_{\mathrm{T}}$ measured at forward rapidity in pp collisions at $\sqrt{s}$ = 13 TeV by the LHCb Collaboration [194]. The data are compared to model calculations  [50, 33, 11, 101] (see text for details). Right panel: comparison of the $\Lambda_{\mathrm{b}}^{0}$ production asymmetry measured by the LHCb Collaboration at $\sqrt{s}$ = 7 TeV with PYTHIA 8 predictions [33, 11].
@@ III.4 Measurements as a function of multiplicity
The observation of substantially different charm baryon-to-meson ratios in pp collisions compared to $\mathrm{e^{+}e^{-}}$ collisions evidences differences in the hadronization process in the two systems. Though at high-$p_{\mathrm{T}}$ the ratios seem to approach system-universal values, suggesting the presence of an energy scale delimiting the modification of the hadronization process, the absence of clear differences between the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratios measured at LEP and B-factory energies and the similarity of the trends observed in pp collisions in the charm and beauty sectors makes it extremely difficult to reconcile pp and $\mathrm{e^{+}e^{-}}$ data ascribing the difference to a kinematic effect related to the energy of the initial-parton or of the jet, without implicating different properties of the systems. One is instinctively tempted to relate the observed differences to the typically larger number of quarks and gluons produced in hadronic collisions. The parton density is strictly connected to the “hadronic activity”, which is encoded in the measurable hadron multiplicity. Inevitably, the study of the evolution of the heavy-flavour baryon-to-meson ratios with particle multiplicity was expected to provide important information to better understand the modification of the hadronization process and to model it. As it will be discussed later in the section, it is important to stress that it is not just the number of hadrons produced that differentiates pp and $\mathrm{e^{+}e^{-}}$ collisions (at LEP energies the multiplicity can reach values not far from the average ones of pp collisions at the LHC) but rather the way in which hadrons can be formed. As discussed in Sect. II.2.1 the concurrence of several initial scatterings, i.e. of MPI, allow for more and topologically different colour connections, which can impact the hadronization process. Also the hadronization via coalescence, which requires multiple quark wavefunctions to overlap in position and momentum space, is expected to progressively emerge as hadron-formation process with increasing number of quarks produced in the collision. Moreover, given the enhancement of multi-strange hadrons observed in high-multiplicity pp collisions  [199], one naturally wonders whether the production of heavy-flavour hadrons containing a strange quark [199] also increases as particle multiplicity increases and whether, if present, this increase should related to heavy-flavour baryon production.
@@
The ALICE and LHCb experiments measured the evolution of the $\mathrm{D_{s}^{+}}/\mathrm{D^{0}}$, $\mathrm{B^{0}_{s}}/\mathrm{B^{0}}$, $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$, and $\mathrm{\Lambda_{\mathrm{b}}^{0}}/B^{0}$ yield ratios from very low to high charged-particle multiplicity in pp collisions at $\sqrt{s}=13$ TeV to provide additional experimental constraints on the hadronization mechanisms in a high density partonic environment [200, 201].
@@
The measured charm-hadron ratios in Fig. 18 for the lowest and highest multiplicity class are compared to model predictions from MC generators and a statistical hadronization model. The simulations with the PYTHIA 8 event generator were performed with the Monash and the CR-BLC tunes. The measured $\mathrm{D_{s}^{+}}/\mathrm{D^{0}}$ ratios at low and high multiplicity are compatible with PYTHIA 8 expectations both with Monash and CR-BLC tunes. As expected by the $\Lambda_{\mathrm{c}}^{+}$/$\mathrm{D^{0}}$ ratio results shown in Fig. 13, the Monash tune does not reproduce the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio in either multiplicity interval, nor does it show a strong multiplicity dependence. By contrast, the CR-BLC tunes describe the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ decreasing trend with $p_{\mathrm{T}}$, and are closer to the overall magnitude in both multiplicity ranges. All CR-BLC tunes show a clear dependence with multiplicity, qualitatively reproducing the trend observed in data, but the evolution predicted by the three “modes” is different, indicating that the data can be useful for tuning further the model.
@@
The predictions of a canonical-ensemble statistical hadronization (CE-SH) model [202] describe the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratios, reproducing their multiplicity dependence. In the calculation, the underlying charm-baryon spectrum in the calculations is augmented to include additional excited baryon states predicted by RQM. The CE-SH model explains the multiplicity dependence as deriving from the reduced volume size of the formalism towards smaller multiplicity, where a decrease of the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio is a consequence of the strict baryon-number conservation. Such behaviour is also expected for charm-strange mesons relative to non-strange charm mesons because of strangeness-number conservation. However, the resulting prediction reproduces the $\mathrm{D_{s}^{+}}/\mathrm{D^{0}}$ ratio only at low multiplicity, while it overestimates it in the highest multiplicity interval. For the $\mathrm{D_{s}^{+}}/\mathrm{D^{0}}$ ratio only the PDG case is shown since the additional RQM states do not modify significantly the D-meson relative yields with respect to the PDG set.
The LHCb Collaboration measured the production rate of $\mathrm{B^{0}_{s}}$ mesons relative to $\mathrm{B^{0}}$ mesons in pp collisions at a centre-of-mass energy $\sqrt{s}$ = 13 TeV over the forward rapidity interval 2 $<y<$ 4.5 as a function of the event charged particle multiplicity [201]. The multiplicity estimators used in the analysis are the total number of charged tracks reconstructed in the VELO detector ($\mathrm{N_{VELO}}$ tracks), and the subset of VELO tracks that point in the backward direction, away from the LHCb spectrometer ($\mathrm{N_{back}}$ tracks). The results reported in Fig. 19 show that the rate of $\mathrm{B^{0}_{s}}$ production relative to $\mathrm{B^{0}}$ might be enhanced with respect to the one measured in $\mathrm{e^{+}e^{-}}$. The enhancement is observed when charged-particle multiplicity is measured in the full VELO detector, while it is not observed when the charged-particle multiplicity is based on $\mathrm{N_{back}}$ tracks. In collisions with relatively low charged-particle multiplicity, and for B mesons with $p_{\mathrm{T}}$ $>$ 6 GeV$/c$, the rate of $\mathrm{B^{0}_{s}}$ production relative to $\mathrm{B^{0}}$ production is consistent with that measured in $\mathrm{e^{+}e^{-}}$ collisions. The ratio of cross sections calculated by the PYTHIA event generator with and without colour reconnection is compared with the experimental measurements for the $p_{\mathrm{T}}$ interval 0–6 GeV$/c$ in the right panel of Fig. 19. Both sets of PYTHIA calculations show a rise with multiplicity, which is more pronounced when colour reconnection is included. Note that the source of this rise in $\mathrm{B^{0}_{s}}/\mathrm{B^{0}}$ with respect to multiplicity in PYTHIA is not due to a explicit strangeness enhancement model (e.g. Rope hadronization [47] or close-packing [48]) but instead is already an existing consequence of the standard Lund fragmentation modelling, however the exact source of this multiplicity dependence has not yet been studied. Both PYTHIA models reproduce the data within uncertainties though both scenarios give values that are lower than the central values of the data at high multiplicity, possibly indicating a different slope of the rise of the $\mathrm{B^{0}_{s}}/\mathrm{B^{0}}$ ratio with multiplicity. The enhancement of the rate of $\mathrm{B^{0}_{s}}$ production relative to $\mathrm{B^{0}}$ production might be a consequence of a bias arising from the coinciding rapidity regions between the multiplicity estimator and the measurement of interest or it might signal that the effect seen is related to the “local” multiplicity of particles surrounding the B mesons. It would be interesting to further investigate these possibilities with Run 3 data and to perform similar measurements at midrapidity. On the other hand, the results could show evidence that the enhancement is present in collisions with high charged-particle multiplicity, while for relatively low charged-particle multiplicity, and for B mesons with $p_{\mathrm{T}}>$ 6 GeV$/c$, the rate of $\mathrm{B^{0}_{s}}$ production relative to $\mathrm{B^{0}}$ production is consistent with what is measured in $\mathrm{e^{+}e^{-}}$ collisions.

### v3

@@ III.3 Proton–proton collisions
The low-$p_{\mathrm{T}}$ charm baryon enhancement is reflected in the $p_{\mathrm{T}}$-integrated yields and their ratios to the $\mathrm{D^{0}}$ meson. The left panel of Fig. 15 shows for pp collisions at $\sqrt{s}=5.02~{}\mathrm{TeV}$ the $p_{\mathrm{T}}$-integrated production cross sections at midrapidity per unit of rapidity for various open- and hidden-charm mesons ($\mathrm{D^{+}}$, $\mathrm{D_{s}^{+}}$, $\mathrm{D^{*+}}$, and $\mathrm{J}/\psi$[201, 175, 202] and baryons ($\Lambda_{\mathrm{c}}^{+}$ and $\Xi_{\mathrm{c}}^{0}$[15, 20], taking the average of particles and antiparticles, and normalizing to the $\mathrm{D^{0}}$ meson. The experimental data is compared with results of pp simulations done with PYTHIA 8 using both the Monash tune [11] and the tunes implementing CR-BLC [51]. For the open charm meson ratios, the PYTHIA 8 predictions obtained with the different tunes are fairly similar and describe the data within uncertainty. The value of the $\mathrm{D^{*+}}/\mathrm{D^{0}}$ ratio obtained with Monash2013 setting $StringFlav:mesonCvector=1.5$ is also reported. As anticipated in Sect. II.1, values of this parameter around 1.5 or larger are preferable to better describe the data, as shown in the right panel of the same figure for the $p_{\mathrm{T}}$-differential $\mathrm{D^{*+}}$/$\mathrm{D^{0}}$ ratio. However, this is not an artefact of the modelling of $\mathrm{D^{*+}}$ pp collision environments. A discrepancy is existent also when considering $\mathrm{e^{+}e^{-}}$ collisions data from LEP and is a consequence of the parameters StringFlav:mesonCvector and correspondingly StringZ:rFactC being mistuned. The default value of 0.88 for StringFlav:mesonCvector should instead be closer to $\sim$1.5. Contrary to the meson case, and as expected from what discussed about the low-$p_{\mathrm{T}}$ trend of the baryon-to-meson ratios shown in Fig. 13, significant differences between the model expectations and data are observed when considering the baryon-to-$\mathrm{D^{0}}$ ratios. The Monash 2013 tune is observed to underestimate the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ and $\Xi_{\mathrm{c}}^{0}/\mathrm{D^{0}}$ ratios by nearly $8\,\sigma$ and $2.3\,\sigma$, respectively. Comparatively, the CR tunes provide larger baryon-to-meson ratios with the Mode 2 tune well reproducing the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio. Nonetheless, they still underestimate the $\Xi_{\mathrm{c}}^{0}/\mathrm{D^{0}}$ ratio by about $2\,\sigma$, which again is unsurprising when looking at the $p_{\mathrm{T}}$ distributions in Fig. 13. For the $\mathrm{J}/\psi/\mathrm{D^{0}}$ ratio the CR tunes provide a better description than the Monash tune, however, all PYTHIA 8 tunes underestimate the measurement. One should note that in the simulations, as in the experimental measurement, the $\mathrm{J}/\psi$ cross section consists of the prompt and beauty feed-down contributions. From these results, one can conclude that the high baryon-to-meson ratios ($\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$, $\Sigma_{\mathrm{c}}^{0,+,++}$/$\mathrm{D^{0}}$, $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$, and $\Omega_{\mathrm{c}}^{0}$/$\mathrm{D^{0}}$), with magnitudes much larger than those of FF ratio measured in $\mathrm{e^{+}e^{-}}$ collisions, clearly indicate that the hadronisation process is not universal and that modelling it solely with FF and FFunc determined from $\mathrm{e^{+}e^{-}}$ data fails already in pp collisions. The right panel of Fig. 15 shows the same experimental data compared in addition to the SHM+RQM and Catania models. For the SHM+RQM for a hadronisation temperature $T_{\mathrm{h}}=$ 170 MeV. Similarly to what is observed for the $p_{\mathrm{T}}$ differential ratios, the SHM+RQM model describes the $\Lambda_{\mathrm{c}}^{+}$ measurements but underestimates the $\Xi_{\mathrm{c}}^{0}$ one. On the other hand, the Catania model overpredicts the production of the $\Lambda_{\mathrm{c}}^{+}$ baryon but captures the production of the $\Xi_{\mathrm{c}}^{0}$ baryon. For the Catania model the $\mathrm{D^{*+}}/\mathrm{D^{0}}$ value shown is the one for the mesons coming from coalescence, because the fragmentation functions employed are inclusive and give only the total contribution to D meson including the $\mathrm{D^{*+}}$ feed-down. The significantly large $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ is due to the combination of both a lower $\mathrm{D^{0}}$ and a larger $\Lambda_{\mathrm{c}}^{+}$ yield with respect to experimental data. However, it has to be noticed that the value in the plot is the original prediction employing the width corresponding to the mean square radii of hadrons from the quark model, hence without any free parameter; a moderate modification of the width parameter would allow to get quite close to the data.
@@
Figure 17: Left panel: ratios $\mathrm{\Lambda_{\mathrm{b}}^{0}}/\mathrm{B^{0}}$ as a function of $p_{\mathrm{T}}$ measured at forward rapidity in pp collisions at $\sqrt{s}$ = 13 TeV by the LHCb Collaboration [14]. The data are compared to model calculations  [69, 51, 11, 116] (see text for details). Right panel: comparison of the $\Lambda_{\mathrm{b}}^{0}$ production asymmetry measured by the LHCb Collaboration at $\sqrt{s}$ = 7 TeV [208] with PYTHIA 8 predictions [51, 11].
@@ III.4 Measurements as a function of multiplicity
The observation of substantially different charm baryon-to-meson ratios in pp collisions compared to $\mathrm{e^{+}e^{-}}$ collisions evidences differences in the hadronization process in the two systems. Though at high-$p_{\mathrm{T}}$ the ratios seem to approach system-universal values, suggesting the presence of an energy scale delimiting the modification of the hadronization process, the absence of clear differences between the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratios measured at LEP and B-factory energies and the similarity of the trends observed in pp collisions in the charm and beauty sectors makes it extremely difficult to reconcile pp and $\mathrm{e^{+}e^{-}}$ data ascribing the difference to a kinematic effect related to the energy of the initial-parton or of the jet, without implicating different properties of the systems. One is instinctively tempted to relate the observed differences to the typically larger number of quarks and gluons produced in hadronic collisions. The parton density is strictly connected to the “hadronic activity”, which is encoded in the measurable hadron multiplicity. Inevitably, the study of the evolution of the heavy-flavour baryon-to-meson ratios with particle multiplicity was expected to provide important information to better understand the modification of the hadronization process and to model it. As it will be discussed later in the section, it is important to stress that it is not just the number of hadrons produced that differentiates pp and $\mathrm{e^{+}e^{-}}$ collisions, but rather the way in which hadrons can be formed. It is interesting to compare the multiplicities of charged particles measured in different collision systems and at different energies. A compilation of average charged particle multiplicities measured in hadronic and leptonic collisions is reported in Fig. 19.6 of Ref. [68]. A precise comparison should take into account and correct for the possible inhomogeneities in the criteria adopted to define the set of particles that are counted, like possible different upper limits on particles lifetimes, kinematic cuts not corrected for, event classes (e.g. inelastic or non-single diffractive for hadronic collisions). Implementing these corrections is far beyond the goal of the qualitative comparison that we want to make here. Values of average multiplicity around 28, with quite large variances [210], are reached at top LEP energies. This number is close to the multiplicities of charged particles with $|eta|<2.5$ in pp collisions at LHC energies (2.36 - 13 $\mathrm{TeV}$) and to the total one in $\mathrm{p\overline{p}}$ collisions at $\sqrt{s}\approx 540$ $\mathrm{GeV}$ measured by UA5 experiment [211]. In pp collisions at LHC, measurements of the total average multiplicities are not available: without aiming at a precise estimate, extrapolations to infinite pseudorapidity range can be done by using either UA5 data [212] or pythia to estimate the ratio of the multiplicity value in a given finite $\eta$ interval ($|eta|<0.5$ or $|eta|<2.5$) to the multiplicity obtained without $\eta$ cuts. In this way, total average multiplicities in pp collisions ranging from about 60 to about 85 from $\sqrt{s}=2.36$ $\mathrm{TeV}$ to $\sqrt{s}=13$ $\mathrm{TeV}$ can be estimated. Therefore, the total average multiplicities at LEP are just a factor 2-3 smaller than those at the LHC. A similar factor applies between the average multiplicities at LEP and at B factories where, as discussed in Sect. III.1, consistent values of $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio are found. Given that at LEP a considerable fraction of particles is part of high-energy jets (most charged particles are within ±$\pm 3$ units in rapidity calculated w.r.t. the sphericity axis [213]), ”local” multiplicities in finite $\eta$ intervals are actually comparable to those of LHC. These considerations should make it clear that it is unlikely that the differences among the relative charm-hadron abundances observed in $\mathrm{e^{+}e^{-}}$ and pp collisions can be related to a simple increase of parton and charged-particle multiplicity without considering the different ways by which the parton and hadrons are formed. The connection of charged-particle multiplicity with MPI in hadronic collisions is a fundamental motivation for studying how charm-quark hadronisation evolves with multiplicity in these collisions. As discussed in Sect. II.2.1, the concurrence of several initial scatterings, i.e. of MPI, allow for more and topologically different colour connections, which can impact the hadronization process. Also the hadronization via coalescence, which requires multiple quark wavefunctions to overlap in position and momentum space, is expected to progressively emerge as hadron-formation process with increasing number of quarks produced in the collision. Moreover, given the enhancement of multi-strange hadrons observed in high-multiplicity pp collisions  [214], one naturally wonders whether the production of heavy-flavour hadrons containing a strange quark [214] also increases as particle multiplicity increases and whether, if present, this increase should be related to heavy-flavour baryon production.
@@
The evolution of the $\mathrm{D_{s}^{+}}/\mathrm{D^{0}}$, $\mathrm{B^{0}_{s}}/\mathrm{B^{0}}$, $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$, and $\Lambda_{\mathrm{b}}^{0}/{\mathrm{B}^{0}}$ cross-section ratios with charged-particle multiplicity was studied in pp collisions at $\sqrt{s}=13$ TeV. Figure 18 shows ALICE measurements of $\mathrm{D_{s}^{+}}/\mathrm{D^{0}}$ and $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratios at low and high multiplicities [23], while the $\mathrm{B^{0}_{s}}/\mathrm{B^{0}}$ and $\Lambda_{\mathrm{b}}^{0}/{\mathrm{B}^{0}}$ trends measured by LHCb [215, 14] are shown in Fig. 19 and Fig. 20, respectively. As discussed in what follows, these results provide important experimental constraints on the modification of the hadronization process from low to high density partonic environments.
@@
The $\mathrm{D_{s}^{+}}/\mathrm{D^{0}}$ ratio at midrapidity does not change significantly from the lowest (top left panel of Fig. 18) to the highest (top right panel) multiplicity class studied by ALICE. Within the uncertainties, it is also independent of $p_{\mathrm{T}}$ in the measured $p_{\mathrm{T}}$ interval and in line with $\mathrm{e^{+}e^{-}}$ values. The $\mathrm{B^{0}_{s}}/\mathrm{B^{0}}$ ratio measured by LHCb at forward rapidity, 2 $<y<$ 4.5, may instead indicate a dependence on multiplicity of the rate of $\mathrm{B^{0}_{s}}$ production relative to $\mathrm{B^{0}}$ one. At high multiplicity, LHCb data favour values higher than those found in $\mathrm{e^{+}e^{-}}$ collisions at LEP, which are compatible and more precise than those obtained at B factories. In collisions with relatively low charged-particle multiplicity, and for B mesons with $p_{\mathrm{T}}$ $>$ 6 GeV$/c$, the $\mathrm{B^{0}_{s}}/\mathrm{B^{0}}$ ratio is consistent with that measured at LEP. In order to properly interpret the difference of the trends observed for $\mathrm{D_{s}^{+}}/\mathrm{D^{0}}$ and $\mathrm{B^{0}_{s}}/\mathrm{B^{0}}$ at low $p_{\mathrm{T}}$, few caveats must be taken into account. First, in the $\mathrm{D_{s}^{+}}/\mathrm{D^{0}}$ case the measurement is limited to $p_{\mathrm{T}}>2$ $\mathrm{GeV}/c$. As mentioned, at $p_{\mathrm{T}}$ larger than the B-meson mass, also the $\mathrm{B^{0}_{s}}/\mathrm{B^{0}}$ ratio does not show a significant dependence on multiplicity. Second, the multiplicity estimators used in the LHCb analysis are the total number of charged tracks reconstructed in the VELO detector ($\mathrm{N_{VELO}}$ tracks), and the subset of VELO tracks that point in the backward direction, away from the LHCb spectrometer ($\mathrm{N_{back}}$ tracks). The $\mathrm{B^{0}_{s}}/\mathrm{B^{0}}$ enhancement is observed when charged-particle multiplicity is measured in the full VELO detector, while it is not observed when the charged-particle multiplicity is based on $\mathrm{N_{back}}$ tracks. The ALICE Collaboration also performed the measurement of the $\mathrm{D_{s}^{+}}/\mathrm{D^{0}}$ ratio, as well as that of the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio discussed later, with two different multiplicity estimators using detectors located at different pseudorapidity regions; one determined by the multiplicity of charged particles at midrapidity, thus overlapping with the pseudorapidity window in which the charm hadrons are reconstructed, and one with a significant $\eta$ gap. To be noted that in the former estimator, which is the one used for the measurements reported in Fig. 18, the contribution of the decay particles of the reconstructed charm hadron is subtracted from the estimated multiplicity. Within uncertainties, ALICE found no differences between the $\mathrm{D_{s}^{+}}/\mathrm{D^{0}}$ ratios measured with the two estimators [23]. Further guidance for interpreting the two measurements is provided by the comparison with PYTHIA expectations. The $\mathrm{D_{s}^{+}}/\mathrm{D^{0}}$ ratios at low and high multiplicity are both described by PYTHIA 8 with Monash tune as well as by tunes with CR-BLC. The differences between the predictions are relatively small and, in the measured $p_{\mathrm{T}}$ interval, all tunes expect a tiny reduction of the $\mathrm{D_{s}^{+}}/\mathrm{D^{0}}$ ratio from low to high multiplicity. The $\mathrm{B^{0}_{s}}/\mathrm{B^{0}}$ ratio calculated by the PYTHIA 8 event generator with and without colour reconnection is compared with the experimental measurements for the $p_{\mathrm{T}}$ interval 0–6 GeV$/c$ in the right panel of Fig. 19. Both sets of PYTHIA calculations show a rise with multiplicity, which is more pronounced when colour reconnection is included. Note that this rise in PYTHIA is not due to an explicit strangeness enhancement mechanism (e.g. Rope hadronization [65] or close-packing [66]) and it is already present with the standard Lund fragmentation model. However, the exact cause of this multiplicity dependence has not yet been studied. Both PYTHIA 8 models reproduce the data within uncertainties though both scenarios give values that are lower than the central values of the data at high multiplicity, possibly indicating a different slope of the rise with multiplicity. At higher $p_{\mathrm{T}}$, both PYTHIA 8 simulations well reproduce the data [215]. The enhancement of the rate of $\mathrm{B^{0}_{s}}$ production relative to $\mathrm{B^{0}}$ production at low $p_{\mathrm{T}}$ might be a consequence of a bias arising from the coinciding rapidity regions between the multiplicity estimator and the measurement of interest or it might signal that the effect seen is related to the “local” multiplicity of particles surrounding the B mesons. It would be interesting to further investigate these possibilities with Run 3 data and to perform similar measurements at midrapidity and to extend the measurement of the $\mathrm{D_{s}^{+}}/\mathrm{D^{0}}$ ratio down to $p_{\mathrm{T}}=0$.
@@
As expected by the $\Lambda_{\mathrm{c}}^{+}$/$\mathrm{D^{0}}$ ratio results shown in Fig. 13, the Monash tune does not reproduce the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio in either multiplicity interval, nor does it show a strong multiplicity dependence. By contrast, the CR-BLC tunes describe the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ decreasing trend with $p_{\mathrm{T}}$, and are closer to the overall magnitude in both multiplicity ranges. All CR-BLC tunes show a clear dependence with multiplicity, qualitatively reproducing the trend observed in data, but the evolution predicted by the three “modes” is different, indicating that the data can be useful for further tuning the model.
@@
The predictions of a canonical-ensemble statistical hadronization (CE-SH) model [216] are also compared to the $\mathrm{D_{s}^{+}}/\mathrm{D^{0}}$ and $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratios. In the calculation, the underlying charm-baryon spectrum is augmented to include additional excited baryon states predicted by RQM. On the contrary, for the $\mathrm{D_{s}^{+}}/\mathrm{D^{0}}$ ratio the prediction obtained considering only the hadron states listed in the PDG is shown since the additional RQM states do not modify significantly the D-meson relative yields with respect to the PDG set. The CE-SH model predicts a dependence on multiplicity for both ratios, deriving from the reduced volume size of the formalism towards smaller multiplicity combined with quantum-number conservation. The resulting modification of the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio reproduces the multiplicity dependence observed in data and interpreted as a decrease of the ratio at low multiplicity as a consequence of the strict baryon-number conservation. However, such behaviour is also expected for charm-strange mesons relative to non-strange charm mesons because of strangeness-number conservation but the resulting prediction reproduces the $\mathrm{D_{s}^{+}}/\mathrm{D^{0}}$ ratio only at low multiplicity, while it overestimates it in the highest multiplicity interval.
It is worth noticing that the measured $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio in the lowest multiplicity class is still higher (in the measured $p_{\mathrm{T}}$ range) than the average of corresponding ratios measured in $\mathrm{e^{+}e^{-}}$ collisions. As it will be further discussed in Sect. III.6, exploring lower multiplicities and extending the measurements to $p_{\mathrm{T}}$ =0 would be of fundamental importance to verify whether the ratio approaches the value measured in $\mathrm{e^{+}e^{-}}$ collisions, as the $p_{\mathrm{T}}$-integrated $\Lambda_{\mathrm{b}}^{0}/{\mathrm{B}^{0}}$ ratio at forward rapidity, which will be now discussed, does.